OpenStudy (anonymous):

I've been asked to find all the isormphic classes of abelian groups of order 760. I google searched and found for 720: Z2×Z2×Z2×Z2×Z3×Z3×Z5 Z4×Z2×Z2×Z3×Z3×Z5 Z4×Z4×Z3×Z3×Z5 Z8×Z2×Z3×Z3×Z5 Z16×Z3×Z3×Z5 Z16×Z9×Z5 Z8×Z2×Z9×Z5 Z4×Z4×Z9×Z5 Z4×Z2×Z2×Z9×Z5 Z2×Z×Z2×Z2×Z9×Z5 How would 760 look?

OpenStudy (kinggeorge):

Abelian groups are relatively to describe. The only thing you need to know is the prime factorization of the order. In this case, we have that \(760=2^3\cdot5\cdot19\). Since finite abelian groups can be written as a direct product of cyclic groups, all you need to to do is realize how many ways you can split \(2^3\cdot5\cdot19\) into factors that are NOT coprime.

OpenStudy (anonymous):

ah ok why are there fewer than 720

OpenStudy (kinggeorge):

There are much fewer because the prime powers are smaller. 720 factors as\[720=2^4\cdot3^2\cdot5\]In this factorization you have both a power of 4 and 2, but in the factorization of 760, we only had a power of 3.

OpenStudy (kinggeorge):

A general formula for the number of abelian groups of order\[n=p_1^{e_1}\cdot p_2^{e_2}\cdot...\cdot p_r^{e_r}\]is given by\[f(e_1)\cdot f(e_2)\cdot...\cdot f(e_r)\]where \(f(e_i)\) is defined to be the number of integer partitions of \(e_i\). See this article for a description of integer partitions. http://en.wikipedia.org/wiki/Partition_(number_theory)

OpenStudy (anonymous):

Ok thank you Im have a look

OpenStudy (anonymous):

* im gonna have a look

OpenStudy (kinggeorge):

Oh, and just to be clear, that factorization of \(n\) I gave is the unique prime factorization.

OpenStudy (kinggeorge):

Sorry, I think I confused myself a little bit when writing the groups of order 760. There are actually 3 groups of that order, and they're actually\[\mathbb{Z}_{760}\]\[\mathbb{Z}_2\times\mathbb{Z}_{380}\]\[\mathbb{Z}_2\times\mathbb{Z}_2\times\mathbb{Z}_{190}\]When I said you need to look at the factors that are not coprime, I left out the additional restriction that each factor need to be divisible by the previous factor.

OpenStudy (kinggeorge):

So in particular, if you have an abelian group of order \(p_1^{e_1}\cdot...\cdot p_{r}^{e_r}\), then you will always have at least one subgroup of order \(p_1\cdot...\cdot p_r\).

OpenStudy (anonymous):

interesting.. still a little confused with this stuff but you cleared it up for me thanks!

OpenStudy (kinggeorge):

You're welcome. Feel free to ask me any more questions on this stuff.

OpenStudy (anonymous):

Will do!